One of three angles is 5 degrees more than the smallest angle. the third angle is three times the middle angle. The sum of the three angles is 120 degrees. Find the measure of each angle.
Answer by RAY100(1637)

Let x,,y,,and z be the angles from smallest to largest
.
y= x+5
.
Z= 3y = 3(x+5) = 3x +15
.
.
x+y+z=120,,,,,,subst
.
x +(x+5) +(3x+15) =120,,,,simplify by collecting like terms
.
5x +20 =120,,,,,,subst 20
.
5x =100,,,,,divide by 5
.
x=20,,,,,,,,,y=x+5 =20+5=25,,,,,,z= 3y =3*25 = 75
.
.
check,,,(20) + (25) + 75 = 120,,,,,ok